1.     A 700.0-N sign hangs from the end of a uniform strut. The strut is 4.0 m long and weighs 800.0 N. The strut is supported by a hinge at the wall and by a cable whose other end is tied to the wall at a point 3.0 m above the left end of the strut. Find the tension in the supporting cable and the force of the hinge on the strut.




1
Expert's answer
2021-11-25T16:23:37-0500
"\frac{3}{5}T=600+400\\T=1667\ N"

"F=\frac{4}{5}1667=1333\ N"